Explain the concept of average propensity to consume (APC) and average propensity to save (APS) .

Given the following information:

Consumption Funciton = C = 1600 + 0.4Yd

Investment = RM1200

Government Expenditure = RM 1180

Taxes = RM6

Derive new consumption function and saving function for this economy.

Calculate the national income equilibrium for a three sector economy using AS=AD Approach and Leakage = Injection Approach. 

Draw the curve for the national income equilibrium for a three sector economy using AS=AD Approach and Leakage = Injection Approach. 


1
Expert's answer
2022-02-01T10:14:33-0500

APC and APS concept

Average propensity to consume (APC), is the expenditure on consumption ratio, to disposable income. In other words, it can be known to be the ratio of total consumption to total income. It measures the income percentage spent, other than saved. On the other hand, Average propensity to save (APS), is the ratio of saved amount of income, to total disposable income. Ratio of savings to total income. APS, is however greatly influenced by the APC. The value of APS can be negative at times, at the point when consumer’s expenditure on consumption exceeds total income. APS can also be zero, at autonomous consumption when income of the consumer is zero. (Wang, 2018)

Deriving Consumption and Savings Function

the consumption function has been given as;

"C=1600+0.4Yd"

"I=1200"

"G=1180"

"T=6"

but we know,;

"C=a+bY" and

"Yd=Y-T"

substituting the given values in the unknown part of the equation we get;

"C=1600+0.4(Y-T)"

"C=1600+0.4(Y-6)"

"C=1600+0.4Y-2.4"

we get the new consumption function as;

"\bold{C=1597.6+0.4Y}"


From the consumption function, we obtain the savings function given as;

"S=-a+(1-b)Y"

substituting we get

"S=-1597.6+(1-0.4)Y"

"\bold{S=-1597.6+0.6Y}"


Calculating national income in a three sector economy

At equilibrium,

"AS=AD"

"Injections=Leakages"

for a three sector economy, "Y=C+I+G-T" (Tihtina, 2021)

"Y=AE"

"Y=injections, AE=leakages"

"Y=C+S+T"

"AE=C+I+G"

"C+S+T=C+I+G"

cancelling C out we get,

"S+T=I+G" equivalent to "Leakages = Injections" at equilibrium.

"(-1597.6+0.6Y)+6=1200+1180"

"0.6Y=2374+1597.6"

national income at equilibrium is;

"Y=6619.33"


National Equilibrium Curve

At the point E in the graph leakages are equal to injections. Beyond this point, leakages exceed injections and no further production is taken. Before this point injections exceeded leakages allowing for more production. The Point E also marks the nation GDP at Y.





Reference.

Tihtina, A. (2021). Three-Sector Keynesian Model of Business Cycles (Doctoral dissertation, ASTU). http://213.55.101.20:8080/xmlui/handle/123456789/1823

Wang, S. (2018). Consumer Theory. In Microeconomic Theory (pp. 37-67). Springer, Singapore. https://doi.org/10.1007/978-981-13-0041-7_2